# IRIS: An Immersive Robot Interaction System

### v2

@@ Abstract
This paper introduces IRIS, an Immersive Robot Interaction System leveraging Extended Reality (XR). IRIS was designed for immersive robot data collection and interaction across multiple simulators, benchmarks, and real-world scenarios. While existing XR-based data collection systems provide efficient and intuitive solutions for large-scale data collection, they are often challenging to reproduce and reuse. This limitation arises because current systems are highly tailored to simulator-specific use cases and environments. IRIS is a novel, easily extensible framework that already supports multiple simulators, benchmarks, and even headsets. Furthermore, IRIS is able to include additional information from real-world sensors, such as point clouds captured through depth cameras. An unified scene specification is generated directly from simulators or real-world sensors and transmitted to XR headsets, creating identical scenes in XR. This specification allows IRIS to support any of the objects, assets, and robots provided by the simulators. In addition, IRIS introduces shared spatial anchors and a robust communication protocol that links simulations between multiple XR headsets. This feature enables these XR headsets to share a synchronized scene, facilitating collaborative and multi-user data collection. IRIS can be deployed on any device that supports the Unity framework, encompassing the vast majority of commercially available headsets. In this work, IRIS was deployed and tested on the Meta Quest 3 and the HoloLens 2. IRIS showcased its versatility across a wide range of real-world and simulated scenarios, using current popular robot simulators such as MuJoCo, IsaacSim, CoppeliaSim, and Genesis. In addition, a user study evaluates IRIS on a data collection task for the LIBERO benchmark. The study shows that IRIS significantly outperforms the baseline in both objective and subjective metrics. This paper demonstrates the huge potential that IRIS provides for the future of data collection and interaction in the robotics and robot learning community. The code, videos, and documentation can be found from the project link(https://intuitive-robots.github.io/iris-project-page/).
Figure 1: We present IRIS, an Immersive Robot Interaction System designed to support various simulators, benchmarks, and real-world scenarios. The system is highly extensible, allowing seamless integration with new simulators and XR headsets. The images shown were captured using a Meta Quest 3 headset, which leverages the Depth API [1] to enable virtual objects to blend naturally with real-world elements rather than always rendering in the foreground. To differentiate between benchmarks, we use distinct labels for each.
I Introduction
Robot learning relies on diverse and high-quality data to learn complex behaviors [6, 59]. Recent studies highlight that models trained on datasets with greater complexity and variation in the domain tend to generalize more effectively across broader scenarios [34, 51, 19]. By leveraging immersive perspectives and interactions, Extended Reality [^1] Extended Reality is an umbrella term to refer to Augmented Reality, Mixed Reality, and Virtual Reality [5] (XR) is a promising candidate for efficient and intuitive large scale data collection [26, 64] in simulation [26, 64, 46] and real-world scenarios [23, 14]. However, reusing and reproducing current XR approaches for robot data collection for new settings and scenarios is complicated and requires significant effort. This bottleneck arises from three main limitations of current XR data collection and interaction frameworks: asset limitation, simulator limitation, and device limitation. Current approaches suffering from asset limitation [30, 26, 64, 20, 38] can only use predefined robot models and task scenes. Configuring new tasks requires significant effort, since each new object or model must be specifically integrated into the XR application. The vast majority of application are developed for specific simulators or real-world scenarios. This simulator limitation [37, 28, 46, 64] significantly reduces reusability and makes adaptation to new simulation platforms challenging. Additionally, most current XR frameworks are designed for a specific version of a single XR headset, leading to a device limitation [28, 40, 23, 35]. To the best of our knowledge, no existing work has explored the extensibility or transferability of their framework to different headsets. These limitations hamper reproducibility and broader contributions of XR based data collection and interaction to the research community.
In addition to these main limitations, existing XR systems are often designed for single-operator use, prohibiting collaborative data collection. At the same time, controlling multiple robots at once can be very difficult for a single operator, making data collection in multi-robot scenarios particularly challenging [43]. Although there are some works using collaborative data collection in the context of tele-operation [58, 49], there is no XR-based data collection system supporting collaborative data collection. This limitation highlights the need for more advanced XR solutions that can better support multi-robot and multi-user scenarios.
@@ I Introduction
To address all of these issues, we propose IRIS, an Immersive Robot Interaction System. This general system supports various simulators, benchmarks and real-world scenarios. It is easily extensible to new simulators and XR headsets. IRIS achieves generalization across six dimensions: Cross-Scene, Cross-Embodiment, Cross-Simulator, Cross-Reality, Cross-Platform, and Cross-User.
@@
Cross-Scene and Cross-Embodiment allow the system to handle arbitrary objects and robots in the simulation, eliminating restrictions about predefined models in XR applications. IRIS achieves these generalizations by introducing a unified scene specification, representing all objects, including robots, as data structures with meshes, materials, and textures. The unified scene specification is transmitted to the XR application to create and visualize an identical scene. By treating robots as standard objects, the system simplifies XR integration, allowing researchers to work with various robots without special robot-specific configurations. Cross-Simulator ensures compatibility with various simulation engines. IRIS simplifies adaptation by parsing simulated scenes into the unified scene specification, eliminating the need for XR application modifications when switching simulators. New simulators can be integrated by creating a parser to convert their scenes into the unified format. This flexibility is demonstrated by IRIS’ support for Mujoco [57], IsaacSim [36], CoppeliaSim [52], and even the recent Genesis [11] simulator. Cross-Reality enables the system to function seamlessly in both virtual simulations and real-world applications. IRIS enables real-world data collection through camera-based point cloud visualization. Cross-Platform allows for compatibility across various XR devices. Since XR device APIs differ significantly, making a single codebase impractical, IRIS XR application decouples its modules to maximize code reuse. This application, developed by Unity [56], separates scene visualization and interaction, allowing developers to integrate new headsets by reusing the visualization code and only implementing input handling for hand, head, and motion controller tracking. IRIS provides an implementation of the XR application in the Unity framework, allowing for a straightforward deployment to any device that supports Unity. So far, IRIS was successfully deployed to the Meta Quest 3 and HoloLens 2. Finally, the Cross-User ability allows multiple users to interact within a shared scene. IRIS achieves this ability by introducing a protocol to establish the communication between multiple XR headsets and the simulation or real-world scenarios. Additionally, IRIS leverages spatial anchors to support the alignment of virtual scenes from all deployed XR headsets. The Immersive Robot Interaction System makes the following contributions
(1) A unified scene specification that is compatible with multiple robot simulators. It enables various XR headsets to visualize and interact with simulated objects and robots, providing an immersive experience while ensuring straightforward reusability and reproducibility.
(2) A collaborative data collection framework designed for XR environments. The framework facilitates enhanced robot data acquisition.
(3) A user study demonstrating that IRIS significantly improves data collection efficiency and intuitiveness compared to the LIBERO baseline.
@@
TABLE I: Comparison of XR-based system for robots. IRIS is compared with related works in different dimensions.
II Related Work
II-A Teleoperation-Based Data Collection on Real Robots
@@
Collecting data using tele-operation on real robots has been explored by many previous works. Aloha [66] introduced a low-cost teleoperation system that collects real-world demonstrations for imitation learning. A bimanual workspace is set up, where leader robots are used to control the follower robots. Followup work [6] improved the performance, ergonomics, and robustness compared to the original design. In addition, a mobile version of Aloha [18] improved data collection outside of lab settings. GELLO [63] supports a variety robot arms through a 3D-printed low-cost leader robots with off-the-shelf motors. In order to tele-operate dexterous end effectors prior work has retrieved hand motion data through visual hand tracking [49] or customized gloves [59]. In contrast to IRIS, none of these approaches leverages the immersive advantages of XR.
II-B XR-Based Data Collection in Real World
@@
Common approaches that combine XR-based tele-operation with real-world interactions typically visualize a virtual robot to show the user how human actions are mapped to robot actuation [49]. For instance, recent work developed mobile apps to allow data collection in augmented reality without the need for XR headsets [16, 60]. However, leveraging XR headsets, allows for more intuitive robot manipulation [64, 40, 26]. Instead of displaying the virtual robot in a third-person view, Cheng et al. [14], Iyer et al. [23] directly provide the first-person camera feed of the real robot to the user. Many other systems [38, 7, 61, 10, 67, 17] visualize the real-world scene in the headset and control the robot arm either with controllers [67] or hand tracking [61]. XR-based data collection for dexterous hands has also been explored. For example, Arunachalam et al. [8] tracks hand motion using camera and retargets it on the real robot hand. Chen et al. [13] controls robot hand and robot arm at the same time. While these approaches do use XR, the robot data collection and interaction is limited to the real world, as no simulators used in the process.
II-C XR-Based Data Collection in Simulation
@@
Collecting data using real robots has many limitations, such as limited scene and object diversity. For gathering demonstrations more efficiently and opening access to numerous virtual 3D assets, some works have explored fully virtual data collection. For instance, DART [46] runs a cloud-based simulation, and users can collect demonstrations in any virtualized environment from any location. Mosbach et al. [37] collects dexterous hand manipulation data with a special glove device in physics simulations. Although Meng et al. [35] also leverages simulators, their virtual scene is a replica of the real scene, thus the flexibility of simulation is not fully exploited.
This discussion reveals that state-of-the-art, XR-based data collection in simulation has barely been explored. Hence, IRIS presents a significant contribution towards immersive robot data collection and interaction. IRIS not only supports almost all popular simulation environments, but also connects seamlessly with robots in the real world, opening up enormous possibilities in human-involved robot learning in both virtual and real-world settings.
III System Overview
This section introduces the features and technical details of IRIS. An overview of its paradigm is shown in Figure 2.
@@ III System Overview
Figure 2: Paradigms of the system architecture in both simulation (left) and real-world (right) environments. The simulation, headsets, and sensors are connected through a Wi-Fi router. In the left image, the simulation updates the scene using the Scene Publisher, broadcasting the information to all XR headsets on the network. A spatial anchor is used to align the virtual scenes across different headsets, allowing users to share a consistent view of the same scene. In the right image, a sensor generates a point cloud, which is sent to the headset. This allows the operator to view the manipulated object in front of the follower robot, enhancing real-time interaction and feedback.
III-A IRIS
III-A1 Node Communication Protocol

### v3

@@ Abstract
This paper introduces IRIS, an Immersive Robot Interaction System leveraging Extended Reality (XR). Existing XR-based systems enable efficient data collection but are often challenging to reproduce and reuse due to their specificity to particular robots, objects, simulators, and environments. IRIS addresses these issues by supporting immersive interaction and data collection across diverse simulators and real-world scenarios. It visualizes arbitrary rigid and deformable objects, robots from simulation, and integrates real-time sensor-generated point clouds for real-world applications. Additionally, IRIS enhances collaborative capabilities by enabling multiple users to simultaneously interact within the same virtual scene. Extensive experiments demonstrate that IRIS offers efficient and intuitive data collection in both simulated and real-world settings.
Figure 1: We present IRIS, an Immersive Robot Interaction System designed to support various simulators and real-world scenarios.
Keywords: Human-Robot Interaction, Extended Reality, Imitation Learning
1 Introduction
Robot learning relies on diverse and high-quality data to acquire complex behaviors [1, 2]. Recent studies indicate that models trained on more varied and complex datasets generalize more effectively across diverse scenarios [3, 4, 5]. By providing immersive perspectives and interactions, Extended Reality  [^1] Extended Reality (XR) is an umbrella term encompassing Augmented Reality, Mixed Reality, and Virtual Reality [6]. (XR) has emerged as a promising tool for efficient and intuitive large-scale data collection in both simulation [7, 8, 9] and real-world environments [10, 11]. However, existing XR approaches face significant challenges when reused or reproduced in new scenarios, primarily due to three limitations: asset diversity, platform dependency, and XR device compatibility.
Current approaches [12, 7, 8, 13, 14] rely heavily on predefined sets of objects and robot models, thereby exhibiting limited asset diversity. Furthermore, most methods [15, 16, 9, 8] are specifically tailored to particular simulators or real-world conditions, resulting in substantial platform dependency. This limitation significantly reduces reusability and complicates adaptation to different simulation platforms. Additionally, existing XR frameworks [16, 17, 10, 18] are typically optimized for specific XR headset versions, leading to poor device compatibility. Together, these limitations severely constrain reproducibility and broader adoption of XR-based data collection and robot interaction methodologies within the research community.
@@ 1 Introduction
To address these challenges, we propose IRIS—an Immersive Robot Interaction System, demonstrated in Figure1. IRIS is a general and extensible framework that supports various simulators and real-world environments, with compatibility across different XR headsets. It is designed to generalize robot applications with XR across six key features: Cross-Scene, Cross-Embodiment, Cross-Simulator, Cross-Reality, Cross-Platform, and Cross-User.
@@
Cross-Scene enables XR systems to handle arbitrary simulated objects, removing constraints from predefined models. IRIS introduces a unified scene specification representing all objects as data structures with meshes, materials, and textures. This specification is transmitted to XR headsets for consistent scene rendering, with dynamic updates during simulation. Through its flexible and dynamic architecture, IRIS is also the first XR-based system that supports deformable objects manipulation. Cross-Embodiment is achieved by modeling robots as compositions of standard objects, enabling seamless compatibility with diverse robot embodiments without requiring specialized configurations. Cross-Simulator ensures compatibility with a range of simulation engines. Since the unified scene specification is simulator-agnostic, new simulators can be supported by implementing a parser to translate their scenes into this format. This flexibility is demonstrated by IRIS’s support for MuJoCo [19], IsaacSim [20], CoppeliaSim [21], and Genesis [22]. Cross-Reality allows IRIS to operate across both simulated and real-world environments. For real-world applications, IRIS incorporates point cloud visualization using camera data, facilitating immersive data collection. Cross-Platform ensures compatibility across XR devices. IRIS implements its XR application using the Unity framework [23], with modular design separating visualization and interaction logic. This allows developers to deploy the system on new XR headsets by reusing visualization modules and implementing device-specific input handling. IRIS has been successfully deployed on the Meta Quest 3 and HoloLens 2. Cross-User supports collaborative multi-user interaction within a shared scene via a communication protocol that synchronizes XR headsets. This enables coordinated tasks and collective data collection in both virtual and real environments. Table 1 highlights the advantages of IRIS over existing XR-based systems across these features.
The contributions of IRIS are summarized as follows: (1) A unified scene specification that integrates seamlessly with multiple robot simulators, enabling consistent visualization and interaction across diverse XR headsets, while promoting reproducibility and reusability. (2) The first XR-based system to support deformable object manipulation in robot simulators, allowing realistic interaction and data collection for soft-body tasks. (3) A collaborative, multi-user framework for XR applications that enhances robot data collection through synchronized interactions in shared virtual or physical environments.
2 Related Work
@@
Table 1: Comparison of XR-based system. IRIS is compared with related works in seven aspects.
@@
Teleoperation-Based Data Collection on Real Robots. Collecting data using tele-operation on real robots has been explored by many previous works. Aloha [30] introduced a low-cost teleoperation system that collects real-world demonstrations for imitation learning. A bimanual workspace is set up, where leader robots are used to control the follower robots. Followup work [1] improved the performance, ergonomics, and robustness compared to the original design. In addition, a mobile version of Aloha [31] improved data collection outside of lab settings. GELLO [32] supports a variety of robot arms through a 3D-printed low-cost leader robots with off-the-shelf motors. In order to tele-operate dexterous end effectors, prior work has retrieved hand motion data through visual hand tracking [33] or customized gloves [2]. In contrast to IRIS, none of these approaches leverages the immersive advantages of XR.
@@
XR-Based Data Collection in Real World. Common XR systems show virtual robots to help users understand how their movements control real robots [33]. For instance, recent works developed mobile apps to allow data collection in augmented reality without the need for XR headsets [34, 35], which allow more intuitive robot manipulation [8, 17, 7]. Instead of displaying the virtual robot in a third-person view, Cheng et al. [11], Iyer et al. [10] directly provide the first-person camera feed of the real robot to the user. Other systems [14, 36, 37, 28, 25, 24] visualize the real-world scene in the headset and control robot arms with controllers [25] or hand tracking [37]. XR-based data collection for dexterous hands has also been explored. For example, Arunachalam et al. [38] tracks hand motion using camera and retargets it on the real robot hand. Chen et al. [39] controls robot hand and robot arm at the same time. While these approaches do use XR, the robot data collection and interaction is limited to the real world, with no simulators used in the process.
@@
XR-Based Data Collection in Simulation. Real robot data collection is limited by available environments and objects. Virtual data collection offers a more efficient way to gather demonstrations while providing access to extensive 3D asset libraries. For instance, DART [9] runs a cloud-based simulation, and users collect demonstrations in any virtualized environment from any location. Mosbach et al. [15] collects dexterous hand manipulation data with a special glove device in physics simulations. Although Meng et al. [18] also leverages simulators, their virtual scene is a replica of the real scene, thus the flexibility of simulation is not fully exploited.
3 System Overview
This section presents the hardware and software architecture of IRIS, along with several applications which have been explored in this paper. An overview of its paradigm is shown in Figure 2.
3.1 System Architecture
@@
Figure 2: Paradigms of the system architecture in both simulation (left) and real world (right). All the devices are connected through a Wi-Fi router. In the left image, the simulation updates the scene to all headsets using the SimPublisher. A spatial anchor is used to align the virtual scenes across different headsets. In the right image, a sensor generates a point cloud transmitted to the XR headset, allowing the operator to clearly observe the manipulated object in front of the follower robot.
Node Communication Protocol. The IRIS system operates across simulation or sensor-processing computers and XR headsets. It provides a dedicated Python library, SimPublisher, for use on the computers, in combination with a Unity application deployed on the headsets. Through this library, users are able to render objects remotely on the headsets and manage groups of headsets without the need to write C# code. This approach eliminates the requirement for developing separate Unity projects, thereby substantially reducing the cost and complexity associated with integrating robotic systems into VR applications. This architecture requires a robust network connection between all components. While Robot Operating System (ROS) [40] offers a general communication framework and it is possible to connect to Unity and XR development, it is quite heavy to deploy. IRIS was intentionally designed to be decoupled from the ROS framework in order to support a broader range of simulators and avoid introducing unnecessary dependencies. It built an lightweight communication protocol based on ZeroMQ (ZMQ) [41], and extended it by auto-node discovery features. This design choice enhances modularity and simplifies integration in diverse environments. Importantly, IRIS remains compatible with ROS through a lightweight and flexible communication interface, allowing seamless integration when needed. To ensure node discovery, the master node broadcasts UDP messages to the broadcast port on the network at a specific frequency. The network is built via Wi-Fi or cable. When a new XR node launches, it listens to the broadcast port and receives messages from the master node. Then it extracts connection details from these messages and establishes a ZMQ connection. This protocol achieves Cross-User ability, ensures reliable communication, automatic reconnection, and smooth recovery from disconnections, making it ideal for dynamic multi-device XR systems.
Unified Scene Representation. To visualize simulation scenes and virtual objects in XR headsets, existing solutions [7, 9, 8, 12] rely on predefined models, which limits flexibility for incorporating new objects and robots. IRIS overcomes this limitation with a Unified Scene Representation (USR) parsed directly from simulations. Unlike URDF, USR can be serialized with meshes and texture information and is adapted to Unity’s GameObject structure, allowing scenes to be reconstructed natively in Unity. Serving as a middleware layer between simulators and Unity, USR enables objects to be generated in Python, serialized, and seamlessly loaded into Unity without additional conversion. The USR includes all objects with their geometry, meshes, materials, and textures. All the objects is loaded in this specification using a kinematic tree structure and serialized into byte format. IRIS application rebuild an identical scene upon received this specification from simulation node. IRIS provides a custom Python library named SimPublisher that automatically generates specifications from simulation data, then it continuously collects simulation states and transmits them to headsets at a fixed frequency. This scene specification enables IRIS to process all kinds of robots and objects in simulation, facilitating both Cross-Scene and Cross-Embodiment capabilities. The USR is a general definition that does not rely on any specific simulator. Hence, IRIS can be easily adapted to various simulators by implementing a new simulation parser to generate a scene specification from the simulator and a new publisher to update the states of the scene. Currently, IRIS supports scene parsers for MuJoCo, IsaacSim, CoppeliaSim, and Genesis, with the potential to be extended to other simulation engines as desired. This demonstrates that IRIS can be easily adapted to various benchmarks and simulators, highlighting its Cross-Simulator capability.
Multiple Headsets Compatibility. IRIS implements an XR application using Unity. This application can be directly deployed to other headset platforms using the Unity deployment pipeline, showcasing IRIS’s Cross-Platform capability. Currently, IRIS has been tested on HoloLens 2 [42] and Meta Quest 3 [43]. Due to Meta Quest 3 visualization resolution is better than HoloLens 2, this paper conducted experiments and displayed XR scenes using this headset.
Intuitive Robot Control Interface. In data collection tasks or robot interaction, robot control interfaces are used to operate the robot in both simulated and real-world environments. Based on prior work [7, 10], IRIS implemented Kinesthetic Teaching (KT) and Motion Controller (MC) as its default robot controllers. The details of these two interfaces are in the Appendix. C. These two methods were used and evaluated with other interfaces in Sec. 2. IRIS’s flexible framework allows users to easily customize and implement additional control interfaces, including hand tracking, gloves, smartphones, and motion tracking systems. Fig. 14 shows how these two interfaces work.
3.2 System Application
IRIS is a versatile platform with significant potential for robot learning and robotics research community. This section outlines several applications that have been explored in this work.
General Manipulation Data Collection. Through its flexible framework design, IRIS supports four simulators and various robot manipulation benchmarks. IRIS has been tested in some MuJoCo-based benchmarks including Meta World [44], LIBERO [45], RoboCasa [46], robosuite [47], Fancy Gym [48], and CoppeliaSim-based benchmark like PyRep [49], Colosseum [50]. Robots can be operated using either our default controllers (Sec. 2) or user-customized controllers.
Deformable Object Manipulation. IRIS supports deformable object by dynamically updating the mesh in real time, making it possible to train and test robotic algorithms for tasks that involve soft objects. As far as we know, no existing work has explored the manipulation of deformable objects using by rendering them in XR headsets. This paper conducted an experiment (Sec. 7) to validate the data collected by IRIS based on IsaacSim.
Figure 3: Collaborative manipulation in simulation via IRIS. The left image shows the collaborative manipulation for handing over a hammer between two Franka Panda robots by KT, and the right image shows that collaborative manipulation for handing over a red board between two Aloha 2 Arms by MC.
Collaborative Manipulation. Collaborative manipulation, where multiple users provide demonstrations simultaneously, is vital for human-robot systems [51]. Previous approaches [33] often use multiple screens, lacking XR’s immersion, and typically limit control to one person while others merely observe [29]. IRIS’s communication protocol enables seamless integration of devices for controlling multiple robots in shared scenes, supporting additional XR headsets with minimal setup for dynamic collaborative environments. Fig. 3 shows collaborative manipulation for handover task.
Figure 4: Playing table tennis with RL agent in Fancy Gym environment. The RL agent policy is trained with Deep Black-Box Reinforcement Learning (BBRL) [52, 53]
High-Dynamic Task Data Collection and Interaction Prior work used keyboards, 3D mice, or smartphones [54, 55, 45, 56], which are limited for complex motions. With the support for motion controllers, IRIS can capture precise user movements. This paper demonstrates the usage of motion controllers in a simulated table tennis task against an RL agent trained with BBRL [52]. Fig. 4 presents this application and an experiment of using collected data to train policies used in Sec. 7.
Figure 5: IRIS Real-world Application. This setup features two Franka robots: a leader robot controlled by a user wearing a Meta Quest 3 headset and a follower robot that mirrors its movements. A depth camera captures the environment for real-time point cloud visualization in XR.
Real World Teleoperation and Data Collection. IRIS not only supports simulation but can also be used in the real-world setting. To minimize physical obstruction from humans, tele-operation using XR is commonly employed for collecting such data. Current approaches [10, 11] that utilize video streaming to XR headsets are restricted to fixed viewpoints. IRIS overcomes this limitation by projecting point cloud from depth cameras to XR headsets, ensuring both immersion and interactivity. Fig. 5 shows the real-world application, and the details are in the Appendix. B.5.2
4 Experiment
In this section, we focus on three questions to demonstrate the effectiveness and scalability of IRIS: (1) How effective and intuitive is the IRIS system for data collection in terms of user experience? (2) Can data collected by IRIS be effectively used for policy training in simulation? (3) Is IRIS suitable for data collection in real-world scenarios? To address these questions, we evaluate the performance of IRIS across three groups of tasks, with detailed experimental settings presented in Appendix D.
4.1 User Experience Evaluation
Interface Task 1 Task 2 Task 3 Task 4
KB (LIBERO) 0.90 0.725 0.750 0.500
3M (LIBERO) 1.00 0.950 0.375 0.900
KT (Ours) 1.00 1.00 0.975 0.950
MC (Ours) 1.00 0.900 0.975 1.00
Table 2: Success rate of four interfaces. KT and MC lead to higher success rate across four tasks.
To assess the efficiency and intuitiveness of IRIS data collection application, a pilot study was conducted by collecting demonstrations for LIBERO benchmark tasks [45]. Four tasks (Fig. 17 in the Appendix) were selected in the dimension of translation, rotation, and compound movement, including close the microwave, turn off the stove, pick up the book in the middle and place it on the cabinet shelf, and turn on the stove and put the frying pan on it. The control interface baselines for this study are two control interfaces from LIBERO: the Keyboard (KB) and the 3D Mouse (3M).
Study Design This study involved eight participants with no prior experience using IRIS or XR headsets. They evaluated each interface through both objective and subjective metrics. Objective measurements included success rate and average time per task, while subjective assessments were gathered via a questionnaire based on the UMUX framework [57]. The questionnaire evaluates each interface by a 7-point Likert scale in four dimensions including Experience, Usefulness, Intuitiveness, and Efficiency. This paper employed the Kruskal-Wallis test [58] for a better and more robust statistical analysis for the study.
Study Result In the result of objective metrics, Tab. 2 and Fig. 6 present the success rates and the average time consumed for each interface across the tasks. The result shows a success rate of over $90\%$ across all four tasks when using the KT and MC interfaces from IRIS, and these XR-based interface significantly outperformed the non-XR interface ($p<0.05$ 222level of statistical significance by Kruskal-Wallis test ) in all conditions except MC in Task 2. The KT and MC interfaces consistently demonstrate lower task completion times than KB and 3M ($p<0.05$) particularly for Task 3 and Task 4, indicating higher efficiency. The result of subjective result is shown in Fig. 6. The KT and MC interfaces consistently receive significant ($p<0.05$) higher scores than KB and 3M across all criteria, indicating positive user perception and ease of use. This study demonstrates that IRIS outperformed baseline interfaces in both objective and subjective metrics, indicating it provides a more intuitive and efficient approach for data collection.
Figure 6: The first row: average task completion time for each interface across tasks. The second row: subjective evaluation scores of four metrics. Our interfaces, KT and MC in XR setting, are faster in data collection with better user experience.
4.2 Policy Evaluation in Simulation
Figure 7: Performance comparison of policies trained on different datasets across LIBERO tasks
General Manipulation To evaluate the quality of data collected using IRIS, we employ two standard imitation learning algorithms: BC-Transformer [59] and BESO [60]. These models are trained separately on datasets collected with IRIS and on the original LIBERO dataset, with results shown in Figure 7. To ensure a fair comparison, we collect the same number of trajectories using IRIS as in LIBERO, using only the MC interface (instead of KT), since LIBERO operates in Cartesian space rather than joint space. Each model is trained for 50 epochs with three random seeds to capture performance variance, and all experiments use identical training parameters. The results demonstrate that the quality of data collected by IRIS is on par with the original LIBERO dataset.
Deformable Objects We also evaluate the data collected by IRIS from deformable object manipulation. Three tasks were designed to evaluate the data including Fold Cloth, Lift Teddy, and Stow Teddy. The policy used in this experiment is the U-Net diffusion model [61]. Observations are robot EEF pose, depth, and image data. The success rate of each task are Fold Cloth: $0.97\pm 0.018$, Lift Teddy: $0.90\pm 0.035$, and Stow Teddy: $0.85\pm 0.053$.
Dynamic Task Data Collection To evaulate the data quality of highly dynamic task collected by IRIS, this paper uses the table tennis from Fancy_Gym [48] by motion controllers. The observation includes bat proprioceptive state and dual camera images, and the action is the desired bat position and orientation in task space. Fig. 8(a) shows the performance of imitation learning models [62, 63] trained on the collected data, using ball interception rate and successful return rate as evaluation metrics. These experiments validate the data from three data collection scenarios in Section 3.2. The results demonstrate that IRIS collects data of comparable quality to traditional methods, while offering significantly greater efficiency.
(a) Policy Performance of Table Tennis Task
(b) Policy Performance of Real World Experiment
Figure 8: Performance evaluation of policies trained on IRIS-collected data across diverse scenarios
4.3 Real World Evaluation
This experiment evaluates the effectiveness of IRIS for real-world data collection through two tasks: Cup Inserting and Picking Up Lego. IRIS was compared against Tele-Op, a widely used method for real robot data collection. In Tele-Op [59, 11, 10], users observe the scene via a monitor or from a distance, which is less comfortable and narrows the operator’s field of view. For each task, 30 successful demonstrations were collected from both methods. These two datasets were used to train two BC-Transformer policies with identical hyperparameters. The metrics include data collection success rate (the percentage of successful attempts during data collection) and policy success rate after training. The results (Fig. 8 b) show that IRIS achieves a higher data collection success rate, and that policies trained with IRIS exhibit better quality than those trained with Tele-Op.
(a) Soft object manipulation in IsaacSim
(b) Real-world data collection by IRIS(left) and Tele-Op(right)
Figure 9: The experiment of deformable manipulation and real-world data collection
5 Conclusion
In this work, we introduced IRIS, an innovative framework that seamlessly integrates Extended Reality (XR) technologies with robotics data collection. IRIS addresses key challenges in reproducibility and reusability that are common in current XR-based systems. Its flexible, extendable design supports multiple simulators, benchmarks, real-world applications, and multi-user use cases. User studies demonstrate that IRIS outperforms previous data collection methods, positioning it as a promising solution for future data collection pipelines. As an open-source project, IRIS codebase promotes further research and adaptation across diverse use cases and hardware platforms.
Limitations
Based on our development and usage experience, IRIS has three main limitations. First, the visualization in XR headsets is limited to basic RGB textures for materials. IRIS currently supports deformable objects exclusively in IsaacSim, as other simulators either lack robust deformable object capabilities or haven’t implemented them. This paper selected IsaacSim for initial testing due to its superior deformable object simulation. Finally, IRIS has only been tested on Meta Quest 3 and HoloLens 2, and further evaluation on additional devices would be beneficial.
Acknowledgments
The presented work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 448648559. Xinkai Jiang and Xiaogang Jia acknowledge the support from the China Scholarship Council (CSC).
Appendix A IRIS Framework
A.1 Node Communication Protocol